OpenStudy (amy2012):

Esmeralda went to the mall traveling 9 mph and returned home traveling 12 mph. If the total trip took 7 hours, how long did Esmeralda travel at each speed?

OpenStudy (zoovy23):

r u in clas right now

OpenStudy (amy2012):

no?

OpenStudy (brainzonly):

This makes no sense. How many minutes did she take AT the mall?

OpenStudy (amy2012):

i think it is asking how long did it take to get there at the speed of 9pmh and home at the speed of 12mph with a total round trip time of 7 hours

OpenStudy (brainzonly):

ok..........

OpenStudy (brainzonly):

argh! this is so frustrating. I know time=distance times speed, but...it's asking how long it took to get there and throwing in 2 speeds! :(

OpenStudy (brainzonly):

with no distance...

OpenStudy (happysunshine44):

Math Logic: The mall is a 7 hour round trip. My Logic: WHY WOULD ANYONE DRIVE A 7 HOUR ROUND-TRIP TO GO TO THE MALL?????

OpenStudy (brainzonly):

That's what I spent 8 minutes wondering...

OpenStudy (happysunshine44):

lol

OpenStudy (brainzonly):

Personally, I have to drive 2 hrs to the nearest town (96 miles) and have never been in a mall! lol

OpenStudy (mathmale):

Let's stay on topic, please.

OpenStudy (amy2012):

Rate (times) Time (equals) Distance... R*T=D

OpenStudy (brainzonly):

ok, so if there are 2 times, it would be 9+12? rate is 7? idk..

OpenStudy (mathmale):

There is one distance: That to the mall (or, that from the mall). Call it d. Assume that d is measured in miles. Then the time required to drive to the mall is \[t _{1}=\frac{ d }{ 9\frac{ miles }{ hour }}\]

OpenStudy (mathmale):

and the time req'd to return, at 12 mph, is \[t _{2}=\frac{ d }{\frac{ 12~miles }{ hour } }\]

OpenStudy (brainzonly):

>>

OpenStudy (mathmale):

t sub 1 and t sub 2 MUST add up to 7 hours. Write this equation. There is not enough info given to actually solve for t sub 1 and t sub 2, that is, to obtain numerical values. Or is there some bright person out there who can do that?

OpenStudy (mathmale):

Note that t sub 2 MUST equal 7 hours - t sub 1. We don't know distance d, but we do know that the distances traveled are equal. Thus:\[d=(9mph)(t _{1})=(12mph)(t _{2})\]

OpenStudy (amy2012):

still not understanding how to get the answer :(

OpenStudy (brainzonly):

I think I get it.

OpenStudy (mathmale):

we can substitute \[7 hours - (12 mph)-t _{1}\] for t sub 2. This eliminates t sub 2, and we therefore have one equation for t sub 1. Find t sub 1.

OpenStudy (mathmale):

Please finish this solution, checking back with the original equation to ascertain that you have done everything that was called for.

OpenStudy (sshayer):

 correction no need to find 2d find only d

OpenStudy (amy2012):

4hours at 9mph and 3 hrs at 12mph!

OpenStudy (brainzonly):

yay!

OpenStudy (sshayer):

correct.

OpenStudy (mathmale):

So cool! Keep up the good work.